OpenStudy (anonymous):

please help me understand or see if what I did is correct will give medal.....

OpenStudy (anonymous):

The second part of the new coaster is a parabola. Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c. Describe the direction of the parabola and determine the y-intercept and zeros. The safety inspector notes that Ray also needs to plan for a vertical ladder through the center of the coaster's parabolic shape for access to the coaster to perform safety repairs. Find the vertex and the equation for the axis of symmetry of the parabola, showing your work, so Ray can include it in his coaster plan. Create a graph of the polynomial function you created in Question 4.

OpenStudy (anonymous):

@johnweldon1993 or @SmartyPants1111 could you help i can show you what i have gotten but not sure if it is correct or not

OpenStudy (johnweldon1993):

Lets see what ya got! :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i picked 3 and 6 to to create the unique parabola @johnweldon1993

OpenStudy (anonymous):

so (x-3) (x-6) then I foiled which I got x^2-6x-3x+18 then i grouped (x^2-6x)+(-3x+18) then factored x(x-6) +3(-x+6) then i had to get rid of the negative x so i divide by the negative so I got x(x-6) +3(x-6) so then (x+3=0) which is x=-3 (x-6=0)which is x=6 and also (x-6=0)which is x=6 then if this is correct im not sure where to go from here ..

OpenStudy (anonymous):

anyone please help

OpenStudy (anthonyym):

Your zeros should be x=3 and x=-6 if your function is y=(x-3)(x+6)

OpenStudy (math&ing001):

If you wrote the parabola as (x-3)(x-6), then you don't need to look for the zeros. You have them directly \( (x- \3) (x- \6) \). They would be 3 and 6. For the y-intercept, put x=0 and solve for y.

OpenStudy (math&ing001):

This might help : https://www.desmos.com/calculator/7lypwvbqle

OpenStudy (anonymous):

Im not understanding what I did wrong though did i do to many steps @math&ing001

OpenStudy (anonymous):

this is what I got

OpenStudy (math&ing001):

The whole process is messing up your result. The goal from doing all that work is to get the zeros by obtaining a factored result like (x-a)(x-b) which YOU ALREADY HAVE from the beginning. So why go through all that again. It's pointless.

OpenStudy (anonymous):

I guess i just need to go from the beginning ...Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax2 + bx + c.

OpenStudy (anonymous):

.Ray needs help creating the second part of the coaster. Create a unique parabola in the pattern f(x) = ax^2 + bx + c.

OpenStudy (math&ing001):

So you need to need foil it (x-3)(x-6) = x^2-6x-3x+18 = x^2-9x+18

OpenStudy (anonymous):

ok i understand that

OpenStudy (math&ing001):

So what's the next step ?

OpenStudy (anonymous):

find the y intercept

OpenStudy (anonymous):

by setting the funtion to 0 @math&ing001 not really sure

OpenStudy (math&ing001):

Yeah just setting x to zero, then solving for y

OpenStudy (anonymous):

so x^2-9x+18 =0^2-9(0)+18 so F(0)=18

OpenStudy (math&ing001):

Yep, so the y-intercept would be 18

OpenStudy (anonymous):

cool so now....Describe the direction of the parabola and determine the y-intercept and zeros. ok i have the y- intercept

OpenStudy (math&ing001):

Scroll up, you also have the zeros ;-] As for the direction in this case, it just depends on the sign of "a" in ax^2+bx+c :

OpenStudy (anonymous):

ok so x-3=0 x=3 and x-6=0 so x=6 and because the a is positive it goes up?

OpenStudy (math&ing001):

Right !

OpenStudy (anonymous):

awesome.. so to get the vertex is from when i graph so it would be 4.5

OpenStudy (math&ing001):

Yes, but you could also get it by derivating the function and puting it to zero : f'(x)=2x-9=0 => x=9/2=4.5

OpenStudy (anonymous):

oh i did not know that thank you and the axis of symmetry is that this formula x=-b/2a

OpenStudy (math&ing001):

Yep, correct !

OpenStudy (anonymous):

A is equal to 1 and B is equal to 9 would that be equal to 4.5 like the the vertex?

OpenStudy (math&ing001):

Yes !

OpenStudy (anonymous):

Thank you so much i think ive got it all

OpenStudy (math&ing001):

Welcome !